# Dynamic spillovers and investment strategies across artificial intelligence ETFs, artificial intelligence tokens, and green markets

### v1

@@ 3.1 $R^{2}$ decomposition approach
In this study, we utilize the innovative $R^{2}$ decomposition technique to examine the dynamic linkages among AI-based ETFs, Tokens, and green markets over different time periods [16, 17]. This approach efficiently disaggregates relevance into contemporaneous and lagged components while ensuring standardized data processing. By overcoming the limitations of prior research, which predominantly concentrated on contemporaneous spillover effects, this method broadens the scope of investigation [16]. The total connectedness index (TCI) is computed as the average $R^{2}$ across $k$ multivariate linear regressions:
@@
When $NET_{i}>0$ (or $NET_{i}<0$), the market $i$ is identified as a net contributor (or receiver) of shocks. Furthermore, net pairwise directional connectedness (NPDC) is calculated in a similar manner to NET, but on a pairwise level [16].
@@ 3.2.1 Bilateral hedge ratios and portfolio weights
According to Kroner et al. [21], the dynamic hedge ratio is expressed as:
@@
Following Kroner and Ng [22], the optimal bilateral portfolio weights are given by:
@@ 3.2.2 Minimum variance portfolio
The minimum variance portfolio (MVP) forms lowest possible risk for a set of assets [18], which is widely used in portfolio management to construct a low-volatility portfolio. The optimal weights for MVP are given by:
@@ 3.2.3 Minimum correlation portfolio
The minimum correlation portfolio (MCP) method is another widely used portfolio construction approach, introduced by Christoffersen et al. [19]. This method constructs portfolios by minimizing conditional correlation. The portfolio weights under MCP are computed as follows:
@@ 3.2.4 Minimum Connectedness Portfolio
The minimum connectedness portfolio (MCoP) approach constructs a portfolio by minimizing the pairwise connectedness of indices [20]. The asset weights are computed as follows:
@@ 3.2.5 Portfolio analysis
To evaluate portfolio performance, two key metrics are used: the Sharpe ratio (SR) [23] and hedging effectiveness (HE) [24].
@@
The hedging effectiveness (HE) measures the percentage risk reduction achieved by the portfolio compared to investing in a single asset $i$ [24]. It is given by:
@@ 4.1 Averaged spillover effects
As illustrated in Fig. 2, the left panel of the figure presents the heatmaps of the average overall, contemporaneous, and lagged connectness. The $(i,j)$ element represents the average spillover from asset $j$ to asset $i$, while the $(j,i)$ element denotes the average spillover from asset $i$ to asset $j$. The difference between these values reflects the net average spillover from asset $j$ to asset $i$. The three heatmaps suggest that the level of information transmission among different types of AI assets remains relatively moderate in both TO and FROM connectedness, mirroring the weak correlation between AI stocks and AI tokens [13]. However, the spillover effects within the same asset are relatively stronger.
Furthermore, AI ETFs act as major net spillover transmitters to both green bonds and AI tokens, a finding that aligns with the results reported by Yousaf et al. [2] and Huynh et al. [10]. A more detailed examination reveals that ROBT transmits net spillovers to all assets, including IRBO and clean energy. This result is consistent with Zeng et al.’s work [7], which highlights that the NASDAQ CTA Artificial Intelligence & Robotics Index, tracked by ROBT, also transmits risk to clean energy. The spillover dynamics of AI tokens exhibit strong similarities, potentially driven by the herding effect in the artificial intelligence and big data token markets. Investors appear to be influenced by collective behavior, leading to coordinated market movements that align with prevailing market consensus [25].
@@
A closer examination indicates that the overall and contemporaneous average spillover effects exhibit similar patterns. First, within the same asset category, spillover effects between assets tend to be stronger. For instance, the overall spillovers from IRBO to ROBT and from ROBT to IRBO are 62.37 and 61.14, respectively, indicating the strongest risk transmission between these two assets. This result is consistent with their highest positive correlation, as shown in Fig. 1. A similar pattern is observed in average contemporaneous connectness. Second, within the AI token market, FIL and ICP exhibit the strongest mutual spillover transmission. Third, spillover effects between clean energy assets and AI ETFs are also relatively strong, indicating notable risk transmission between these asset classes.
However, a closer examination of Fig. 2 (a-c) reveals distinctions among the averaged overall, contemporaneous, and lagged spillover effects. Specifically, in the overall analysis, clean energy transmits net spillovers to all assets, including IRBO and bonds, except for ROBT. In the contemporaneous period, all assets transmit net spillovers to green bond. In the lagged period, clean energy transmits spillovers to tokens and bonds.
Observing Fig. 2 (d-f), the three bar charts on the right illustrate the averaged spillover effects among different assets, with each panel representing overall (black), contemporaneous (red), and lagged (green) connectedness. Fig. 2 (d) shows that AI ETFs (IRBO and ROBT) and clean energy have the highest TO connectenss, indicating that they are the primary risk transmitters, with most spillover effects occurring contemporaneously rather than with a lag. Fig. 2 (e) presents FROM connectedness, measuring the extent to which each asset receives risk from others. AI ETFs and clean energy continue to exhibit the most significant FROM values, suggesting that they not only transmit risk but also absorb it from other assets. Lagged effects remain relatively small, indicating that risk transmission is predominantly immediate. Fig. 2 (f) represents NET connectedness, calculated as the difference between TO and FROM spillovers. AI ETFs have the highest net spillover values, reinforcing their role as key risk transmitters, while AI tokens and green bond generally exhibit negative net spillover values, confirming their role as risk recipients. Notably, SPGTCED has a positive net spillover value, suggesting that it still transmits a certain level of risk to other assets. All assets display lagged spillover effects, indicating that NET risk transmission may experience slight delays over time. Overall, these charts suggest that AI ETFs and clean energy serve as the primary sources of risk transmission, whereas AI tokens and green bonds mainly act as risk recipients, with NET spillover effects occurring primarily contemporaneously rather than over extended periods. The empirical results are in line with work of Billah [9], Yousaf et al. [3], and Yousaf et al. [2].
@@ 4.2 Dynamic total connectedness measures
Overall, the strong synchronization between overall and contemporaneous connectedness suggests that risk spillovers primarily occur in real-time rather than with a lag. The consistently low lagged connectedness indicates that market information transmission does not experience significant delays.
@@ 4.3 NET total directional connectedness
Fig. 4 illustrates the time-varying directional connectedness of AI and green assets. From the figure, it can be observed that the overall spillover effects and lagged spillover effects exhibit a closely aligned trend across various asset classes. In contrast, the net contemporaneous spillover effects show relatively weaker fluctuations, with certain assets experiencing significant increases in spillover effects during specific periods.
@@
The AI ETFs and clean energy assets exhibit a similar trend, both reaching a peak around October 2022, followed by a decline, and then an upward trajectory starting in 2024. Their overall spillovers remain predominantly positive throughout the period, consistently serving as the strongest contributors to systemic risk, highlighting their role as key drivers of market fluctuations. Ref. [7] and this study both conclude that clean energy acts as a net transmitter. Additionally, while Zeng et al. identify the AI index as a net transmitter [7], this study finds that ROBT tracking the AI index also serves as a net transmitter. Consistent with Billah’s findings [9], AI tokens likewise act as recipients of shocks within the system.
In contrast, AI tokens and SPGB maintain relatively low levels of overall spillover dynamics, which remain negative for most of the period, except between mid-2023 and mid-2024. Their behavior is almost the opposite of AI ETFs and clean energy assets, indicating that they generally act as recipients of shocks within the system. This is particularly evident during periods of heightened market volatility, where their net spillover inflows become more pronounced. However, it is important to note that the spillover roles of these assets evolve dynamically over time, and in certain periods, they may also serve as risk transmitters.
Overall, Fig. 4 reflects the complex interrelationships between the AI and green asset markets. AI ETFs and clean energy assets exhibit strong similarities, whereas AI tokens and the green bond market follow distinct spillover effect patterns. This suggests that different types of AI and green assets interact differently with systemic risk, highlighting the necessity of further exploration into their interconnectedness under varying market conditions.
@@ 4.4 Net pairwise directional connectedness measures
Fig. 5 illustrates the net pairwise directional spillover. The dynamics of risk transmission from these two ETFs to other assets are highly similar, and ROBT primarily transmits risk to IRBO. Aligning with the findings of Huynh et al. [10], AI funds exhibits high overall spillovers, which is consistent with Ref.  [2]. Moreover, the spillover effects between AI ETFs and AI tokens exhibit a similar pattern, generally increasing throughout the sample period until reaching a peak, followed by a decline, reaching their lowest levels between late 2023 and the first quarter of 2024, before rising again.
@@ 4.5 Net pairwise network
Fig. 6 depicts the net pairwise directional spillover networks, filtering edges below 0.05 to highlight major spillover effects. Blue nodes represent transmitters, while yellow nodes denote receivers. In line with the results documented in Ref. [15], our study also identifies a risk spillover effect from clean energy to green bonds. Again, the AI ETFs is the primary shock transmitters to other assets. Connectedness among AI tokens are weaker than their links with AI ETFs and green assets. The overall and lagged networks suggests the strong spillovers from AI ETFs and clean energy to other assets. ROBT is the largest transmitter, followed by IRBO, and finally clean energy. Yousaf et al. also report that AI ETFs (tokens) are strong (weakest) net transmitters (recipients) during normal market conditions [2]. AI tokens and green bond receive relatively similar level of shocks from other markets. However, the contemporaneous network indicates that IRBO, FIL, RENDER, and SPGB acts as receivers. Green bond receives most shocks from other markets, followed by IRBO. Further, FET, ICP, and NEAR transmit risk, although the level of shock remains relatively low.
Figure 6: Net pairwise directional spillover networks.
@@ 4.6 Robustness test
As shown in Fig. 7, we conduct a robustness analysis to verify the reliability of our findings, employing multiple approaches, including Pearson, Spearman, and Kendall correlation-based $R^{2}$ decompositions, as well as quantile VAR (QVAR) [26]. All TCIs exhibit a largely consistent pattern, reinforcing the robustness of our findings.
@@ 5.1 Bilateral hedge ratios and portfolio weights
      Long/Short       Mean       Std. Dev.       5%       95%       HE       p-value
      IRBO/FET       0.007       0.011       --0.010       0.022       0.018       0.886
      IRBO/FIL       0.008       0.017       --0.018       0.041       0.018       0.855
      IRBO/ICP       0.000       0.017       --0.024       0.037       0.014       0.887
      IRBO/NEAR       0.009       0.014       --0.020       0.030       0.018       0.850
      IRBO/RENDER       0.012       0.010       --0.002       0.029       0.020       0.840
@@
      ROBT/FET       0.004       0.011       --0.013       0.022       0.018       0.886
      ROBT/FIL       0.000       0.017       --0.028       0.028       0.017       0.855
      ROBT/ICP       --0.008       0.017       --0.033       0.027       0.012       0.887
      ROBT/NEAR       0.004       0.013       --0.017       0.024       0.018       0.850
      ROBT/RENDER       0.004       0.011       --0.013       0.023       0.017       0.840
@@
      FET/IRBO       0.185       0.285       --0.225       0.772       0.016       0.000
      FET/ROBT       0.106       0.291       --0.316       0.571       0.015       0.000
      FET/SPGB       --0.520       1.425       --3.129       2.175       0.008       0.002
      FET/SPGTCED       --0.104       0.284       --0.570       0.326       0.009       0.000
      FIL/IRBO       0.139       0.277       --0.261       0.632       0.015       0.000
      FIL/ROBT       --0.015       0.283       --0.468       0.390       0.015       0.000
      FIL/SPGB       0.418       0.830       --0.677       1.963       0.016       0.002
      FIL/SPGTCED       --0.027       0.274       --0.466       0.486       0.012       0.000
      ICP/IRBO       0.035       0.302       --0.421       0.713       0.013       0.000
      ICP/ROBT       --0.120       0.306       --0.527       0.543       0.011       0.000
      ICP/SPGB       --0.165       0.715       --1.623       1.093       0.003       0.002
      ICP/SPGTCED       --0.091       0.266       --0.422       0.462       0.009       0.000
      NEAR/IRBO       0.115       0.375       --0.880       0.624       0.018       0.000
      NEAR/ROBT       0.066       0.327       --0.664       0.503       0.017       0.000
      NEAR/SPGB       0.697       0.906       --1.004       2.157       0.015       0.002
      NEAR/SPGTCED       --0.121       0.425       --1.306       0.373       0.012       0.000
      RENDER/IRBO       0.295       0.250       --0.126       0.699       0.018       0.000
      RENDER/ROBT       0.108       0.331       --0.400       0.777       0.015       0.000
      RENDER/SPGB       0.015       0.971       --1.415       1.618       0.009       0.002
      RENDER/SPGTCED       --0.079       0.305       --0.531       0.434       0.013       0.000
@@
      SPGB/FET       --0.002       0.004       --0.009       0.006       0.010       0.886
      SPGB/FIL       0.002       0.004       --0.003       0.009       0.015       0.855
      SPGB/ICP       0.000       0.003       --0.005       0.005       0.007       0.887
      SPGB/NEAR       0.004       0.004       --0.002       0.011       0.016       0.850
      SPGB/RENDER       0.001       0.003       --0.003       0.006       0.011       0.840
@@
      SPGTCED/FET       --0.004       0.010       --0.022       0.010       0.011       0.886
      SPGTCED/FIL       0.000       0.015       --0.025       0.028       0.014       0.855
      SPGTCED/ICP       --0.009       0.016       --0.036       0.020       0.008       0.887
      SPGTCED/NEAR       --0.003       0.014       --0.034       0.018       0.011       0.850
      SPGTCED/RENDER       --0.002       0.011       --0.020       0.016       0.013       0.840
Second, AI tokens provide limited effectiveness when used to hedge AI ETFs. Hedge ratio means are near zero, HE remains low (0.011-0.020), and $p$-values exceed 0.8, indicating poor risk reduction capabilities. When tokens are long and AI ETFs are short, hedge ratios increase. The highest hedge ratio is observed in RENDER/IRBO (0.295), while NEAR/IRBO achieves the highest HE (0.018), still providing minimal risk mitigation. Compared to Ref. [2], the hedge ratio between AI ETFs and tokens is highly similar. However, the hedging effectiveness differ. This discrepancy may stem from differences in the time period analyzed and sample selection.

### v2

@@ 4.1 $R^{2}$ decomposition approach
In this study, we utilize the innovative $R^{2}$ decomposition technique to examine the dynamic linkages among AI-based ETFs, Tokens, and green markets over different time periods (Balli et al., 2023; Naeem et al., 2024). This approach efficiently disaggregates relevance into contemporaneous and lagged components while ensuring standardized data processing. By overcoming the limitations of prior research, which predominantly concentrated on contemporaneous spillover effects, this method broadens the scope of investigation (Balli et al., 2023). The total connectedness index (TCI) is computed as the average $R^{2}$ across $k$ multivariate linear regressions:
@@
When $NET_{i}>0$ (or $NET_{i}<0$), the market $i$ is identified as a net contributor (or receiver) of shocks. Furthermore, net pairwise directional connectedness (NPDC) is calculated in a similar manner to NET, but on a pairwise level (Balli et al., 2023).
@@ 4.2.1 Bilateral hedge ratios and portfolio weights
According to Kroner and Sultan (1993), the dynamic hedge ratio is expressed as:
@@
Following Kroner and Ng (1998), the optimal bilateral portfolio weights are given by:
@@ 4.2.2 Minimum variance portfolio
The minimum variance portfolio (MVP) forms lowest possible risk for a set of assets (Markowitz, 1952), which is widely used in portfolio management to construct a low-volatility portfolio. The optimal weights for MVP are given by:
@@ 4.2.3 Minimum correlation portfolio
The minimum correlation portfolio (MCP) method is another widely used portfolio construction approach, introduced by Christoffersen et al. (2014). This method constructs portfolios by minimizing conditional correlation. The portfolio weights under MCP are computed as follows:
@@ 4.2.4 Minimum connectedness portfolio
The minimum connectedness portfolio (MCoP) approach constructs a portfolio by minimizing the pairwise connectedness of indices (Broadstock et al., 2022). The asset weights are computed as follows:
@@ 4.2.5 Portfolio analysis
To evaluate portfolio performance, two key metrics are used: the Sharpe ratio (SR) (Sharpe, 1966) and hedging effectiveness (HE) (Ederington, 1979).
@@
The hedging effectiveness (HE) measures the percentage risk reduction achieved by the portfolio compared to investing in a single asset $i$ (Ederington, 1979). It is given by:
@@ 5.1 Averaged spillover effects
Fig. 2 (a-c) present the heatmaps of the average overall, contemporaneous, and lagged connectness. Darker colors in the figure indicate higher levels of risk spillover. The $(i,j)$ element represents the average spillover from asset $j$ to asset $i$, while the $(j,i)$ element denotes the average spillover from asset $i$ to asset $j$. The difference between these values reflects the net average spillover from asset $j$ to asset $i$.
A detailed inspection of Fig. 2 (a-b) indicates that the overall and contemporaneous average spillover effects exhibit similar patterns. First, within the same asset category, spillover effects between assets tend to be stronger. For instance, the overall spillovers from ROBT to IRBO and from IRBO to ROBT are 62.37 and 61.14, respectively, indicating the strongest risk transmission between these two assets. This result is consistent with their highest positive correlation, as shown in Fig. 1. Green assets and AI tokens exhibit a similar pattern. Second, the level of information transmission among different types of AI assets remains relatively moderate, mirroring the weak correlation between AI stocks and AI tokens (Jareño and Yousaf, 2023). Third, spillover effects between clean energy assets and AI ETFs are also relatively strong, indicating notable risk transmission between these asset classes. Moreover, within the AI token market, FIL and ICP exhibit the strongest mutual spillover transmission.
Furthermore, AI ETFs act as major net spillover transmitters to both green bonds and AI tokens, a finding that aligns with the results reported by Yousaf et al. (2024c) and Huynh et al. (2020). A more detailed examination reveals that ROBT transmits net spillovers to all assets, including IRBO and clean energy. This result is consistent with work of Zeng et al. (2024), which highlights that the NASDAQ CTA Artificial Intelligence & Robotics Index, tracked by ROBT, also transmits risk to clean energy. Additionally, the spillover dynamics of AI tokens exhibit strong similarities, potentially driven by the herding effect in the artificial intelligence and big data token markets. Investors appear to be influenced by collective behavior, leading to coordinated market movements that align with prevailing market consensus (Xiaoyang Xu, 2024).
@@
However, a closer look at Fig. 2 (a-c) reveals distinctions among the averaged overall, contemporaneous, and lagged spillover effects. Specifically, in the overall analysis, clean energy transmits net spillovers to all assets, except ROBT. In the contemporaneous period, all assets transmit net spillovers to green bond. In the lagged period, clean energy transmits spillovers to tokens and bonds.
Observing Fig. 2 (d-f), the three bar charts on the right illustrate the averaged spillover effects among different assets, with each panel representing overall (black), contemporaneous (red), and lagged (green) connectedness. Fig. 2 (d) displays the TO connectedness, indicating how much each asset contributes to risk spillovers within the system. AI ETFs (IRBO and ROBT) and clean energy have the highest TO connectenss, indicating that they are the primary risk transmitters, with most spillover effects occurring contemporaneously rather than with a lag. Fig. 2 (e) presents FROM connectedness, measuring the extent to which each asset receives risk from others. AI ETFs and clean energy continue to exhibit the most significant FROM values, suggesting that they not only transmit risk but also absorb it from other assets. Lagged effects remain relatively small, indicating that risk transmission is predominantly immediate. Fig. 2 (f) represents NET connectedness, calculated as the difference between TO and FROM spillovers. AI ETFs have the highest net spillover values, reinforcing their role as key risk transmitters, while AI tokens and green bond generally exhibit negative net spillover values, confirming their role as risk recipients. Notably, SPGTCED has a positive net spillover value, suggesting that it still transmits a certain level of risk to other assets.
Overall, Fig. 2 (d-f) suggest that AI ETFs and clean energy serve as the primary sources of risk transmission, whereas AI tokens and green bonds mainly act as risk recipients, with NET spillover effects occurring over extended periods. The empirical results are in line with work of Billah (2025), Yousaf et al. (2024a), and Yousaf et al. (2024c).
@@ 5.2 Dynamic total connectedness measures
In summary, the strong synchronization between overall and contemporaneous connectedness suggests that risk spillovers primarily occur in real-time rather than with a lag. The consistently low lagged connectedness indicates that market information transmission does not experience significant delays.
@@ 5.3 NET total directional connectedness
Fig. 4 illustrates the time-varying directional connectedness of AI and green assets. In line with Fig. 2 (f), it can be observed that the overall spillover effects and lagged spillover effects exhibit a closely aligned trend across various asset classes. In contrast, the net contemporaneous spillover effects show relatively weaker fluctuations, with certain assets experiencing significant increases in spillover effects during specific periods.
@@
Moreover, the AI ETFs and clean energy assets exhibit a similar trend, both reaching a peak around October 2022, followed by a decline, and then an upward trajectory starting in 2024. Their overall spillovers remain predominantly positive throughout the period, consistently serving as the strongest contributors to systemic risk, highlighting their role as key drivers of market fluctuations. Zeng et al. (2024) and we both conclude that clean energy acts as a net transmitter. Additionally, this study finds that AI funds also serves as a net transmitter, which is consistent with Yousaf et al. (2024c). In line with Billah (2025)’s findings, AI tokens likewise act as recipients of shocks within the system.
In addition, AI tokens and SPGB maintain relatively low levels of spillover dynamics, which remain negative for most of the period, except between mid-2023 and mid-2024. Their overall spillover behavior is almost the opposite of AI ETFs and clean energy assets, indicating that they generally act as recipients of shocks within the system. This is particularly evident during periods of heightened market volatility, where their net spillover inflows become more pronounced. However, it is important to note that the spillover roles of these assets evolve dynamically over time, and in certain periods, they may also serve as risk transmitters.
Generally speaking, Fig. 4 reflects the complex interrelationships between the AI and green asset markets. AI ETFs and clean energy assets exhibit strong similarities, whereas AI tokens and the green bond market follow distinct spillover effect patterns. This suggests that different types of AI and green assets interact differently with systemic risk, highlighting the necessity of further exploration into their interconnectedness under varying market conditions.
@@ 5.4 Net pairwise directional connectedness measures
Fig. 5 illustrates the net pairwise directional spillover. The dynamics of risk transmission from these two ETFs to other assets are highly similar, and IRBO primarily receives risk from ROBT. Aligning with the findings of Huynh et al. (2020), AI funds exhibits high overall spillovers. Moreover, the spillover effects between AI ETFs and AI tokens exhibit a similar pattern, generally increasing throughout the sample period until reaching a peak, followed by a decline, reaching their lowest levels between late 2023 and the first quarter of 2024, before rising again.
@@ 5.5 Net pairwise network
Fig. 6 depicts the net pairwise directional spillover networks, filtering edges below 0.05 to highlight major spillover effects. Blue nodes represent transmitters, while yellow nodes denote receivers. The thickness of edges indicates the strength of spillover effects. In line with the results documented by Tiwari et al. (2022), our study also identifies a risk spillover effect from clean energy to green bonds. Again, the AI ETFs are the primary shock transmitters to other assets. Connectedness among AI tokens are weaker than their links with AI ETFs and green assets. The overall and lagged networks suggests the strong spillovers from AI ETFs and clean energy to other assets. ROBT is the largest transmitter, followed by IRBO, and finally clean energy. Yousaf et al. (2024c) also report that AI ETFs (tokens) are strong (weakest) net transmitters (recipients) during normal market conditions. AI tokens and green bond receive relatively similar level of shocks from other markets.
However, the contemporaneous network indicates that IRBO, FIL, RENDER, and SPGB acts as receivers. Green bond receives most shocks from other markets. Further, FET, ICP, and NEAR transmit risk, although the level of shock remains relatively low.
Figure 6: Net pairwise directional spillover networks.The three panels, from left to right, illustrate overall, contemporaneous, and lagged dependencies.
@@ 5.6 Robustness test
As shown in Fig. 7, we conduct a robustness analysis to verify the reliability of our findings, employing multiple approaches, including Pearson, Spearman, and Kendall correlation-based $R^{2}$ decompositions, as well as quantile VAR (QVAR) (Ando et al., 2022). All TCIs exhibit a largely consistent pattern, reinforcing the robustness of our findings.
@@ 6.1 Bilateral hedge ratios and portfolio weights
Mean Std. Dev. 5% 95% HE p-value
IRBO/FET $0.007$ $0.011$ $-0.010$ $0.022$ $0.018$ $0.886$
IRBO/FIL $0.008$ $0.017$ $-0.018$ $0.041$ $0.018$ $0.855$
IRBO/ICP $0.000$ $0.017$ $-0.024$ $0.037$ $0.014$ $0.887$
IRBO/NEAR $0.009$ $0.014$ $-0.020$ $0.030$ $0.018$ $0.850$
IRBO/RENDER $0.012$ $0.010$ $-0.002$ $0.029$ $0.020$ $0.840$
@@
ROBT/FET $0.004$ $0.011$ $-0.013$ $0.022$ $0.018$ $0.886$
ROBT/FIL $0.000$ $0.017$ $-0.028$ $0.028$ $0.017$ $0.855$
ROBT/ICP $-0.008$ $0.017$ $-0.033$ $0.027$ $0.012$ $0.887$
ROBT/NEAR $0.004$ $0.013$ $-0.017$ $0.024$ $0.018$ $0.850$
ROBT/RENDER $0.004$ $0.011$ $-0.013$ $0.023$ $0.017$ $0.840$
@@
FET/IRBO $0.185$ $0.285$ $-0.225$ $0.772$ $0.016$ $0.000$
FET/ROBT $0.106$ $0.291$ $-0.316$ $0.571$ $0.015$ $0.000$
FET/SPGB $-0.520$ $1.425$ $-3.129$ $2.175$ $0.008$ $0.002$
FET/SPGTCED $-0.104$ $0.284$ $-0.570$ $0.326$ $0.009$ $0.000$
FIL/IRBO $0.139$ $0.277$ $-0.261$ $0.632$ $0.015$ $0.000$
FIL/ROBT $-0.015$ $0.283$ $-0.468$ $0.390$ $0.015$ $0.000$
FIL/SPGB $0.418$ $0.830$ $-0.677$ $1.963$ $0.016$ $0.002$
FIL/SPGTCED $-0.027$ $0.274$ $-0.466$ $0.486$ $0.012$ $0.000$
ICP/IRBO $0.035$ $0.302$ $-0.421$ $0.713$ $0.013$ $0.000$
ICP/ROBT $-0.120$ $0.306$ $-0.527$ $0.543$ $0.011$ $0.000$
ICP/SPGB $-0.165$ $0.715$ $-1.623$ $1.093$ $0.003$ $0.002$
ICP/SPGTCED $-0.091$ $0.266$ $-0.422$ $0.462$ $0.009$ $0.000$
NEAR/IRBO $0.115$ $0.375$ $-0.880$ $0.624$ $0.018$ $0.000$
NEAR/ROBT $0.066$ $0.327$ $-0.664$ $0.503$ $0.017$ $0.000$
NEAR/SPGB $0.697$ $0.906$ $-1.004$ $2.157$ $0.015$ $0.002$
NEAR/SPGTCED $-0.121$ $0.425$ $-1.306$ $0.373$ $0.012$ $0.000$
RENDER/IRBO $0.295$ $0.250$ $-0.126$ $0.699$ $0.018$ $0.000$
RENDER/ROBT $0.108$ $0.331$ $-0.400$ $0.777$ $0.015$ $0.000$
RENDER/SPGB $0.015$ $0.971$ $-1.415$ $1.618$ $0.009$ $0.002$
RENDER/SPGTCED $-0.079$ $0.305$ $-0.531$ $0.434$ $0.013$ $0.000$
@@
SPGB/FET $-0.002$ $0.004$ $-0.009$ $0.006$ $0.010$ $0.886$
SPGB/FIL $0.002$ $0.004$ $-0.003$ $0.009$ $0.015$ $0.855$
SPGB/ICP $0.000$ $0.003$ $-0.005$ $0.005$ $0.007$ $0.887$
SPGB/NEAR $0.004$ $0.004$ $-0.002$ $0.011$ $0.016$ $0.850$
SPGB/RENDER $0.001$ $0.003$ $-0.003$ $0.006$ $0.011$ $0.840$
@@
SPGTCED/FET $-0.004$ $0.010$ $-0.022$ $0.010$ $0.011$ $0.886$
SPGTCED/FIL $0.000$ $0.015$ $-0.025$ $0.028$ $0.014$ $0.855$
SPGTCED/ICP $-0.009$ $0.016$ $-0.036$ $0.020$ $0.008$ $0.887$
SPGTCED/NEAR $-0.003$ $0.014$ $-0.034$ $0.018$ $0.011$ $0.850$
SPGTCED/RENDER $-0.002$ $0.011$ $-0.020$ $0.016$ $0.013$ $0.840$
Second, AI tokens provide limited effectiveness when used to hedge AI ETFs. Hedge ratio means are near zero, HE remains low (ranging from 0.012 to 0.020), and $p$-values exceed 0.8, indicating poor risk reduction capabilities. When tokens are long and AI ETFs are short, hedge ratios increase and HE is still low. The highest hedge ratio is observed in RENDER/IRBO (0.295), while NEAR/IRBO and RENDER/IRBO achieve the highest HE (0.018), still providing minimal risk mitigation.